OpenStudy (anonymous):

Evaluate the logarithm. log 6 1/36

OpenStudy (anonymous):

Homezone?

OpenStudy (anonymous):

sorry i was in another window :)

OpenStudy (anonymous):

ok can you help step by step?

OpenStudy (anonymous):

do you mean log_6 (1/36) ?

OpenStudy (anonymous):

I guess how its written up top is how it is

OpenStudy (anonymous):

(1/36)=6^-2

OpenStudy (anonymous):

log_6 (6^-2)=-2

OpenStudy (anonymous):

ok for you ?

OpenStudy (anonymous):

whered u get -2?

OpenStudy (anonymous):

log_6(6)=1 log_6(6^2)=2 log_6(6^10)=10 log_6(6^-10)=-10

OpenStudy (anonymous):

just as examples

OpenStudy (anonymous):

you can rewrite 1/36 as (1/(6^2)) and then again as 6^-2

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

you don't get it ?

OpenStudy (anonymous):

not really also whered u get =-2

OpenStudy (anonymous):

do you get log_6(6)=1 ?

OpenStudy (anonymous):

you can take the 6' logarithm to 6 once ergo log_6(6)=1

OpenStudy (anonymous):

not really I don't really even know what log means in an equation

OpenStudy (anonymous):

Log is a function, i'd advice you to read up on it.

OpenStudy (anonymous):

if you can't even do 1/36=1/6^2=6^-2. you shouldn't do functions yet

OpenStudy (anonymous):

ok well I need to do this I'm at log 6 (6^-2)

OpenStudy (anonymous):

what do I do next?

OpenStudy (anonymous):

Hi James can you help and try to explain?

OpenStudy (anonymous):

Can you please help?

OpenStudy (anonymous):

Jdoe?

OpenStudy (jdoe0001):

do you know what a negative exponent means?

OpenStudy (anonymous):

no I need help with the problem at the very top

OpenStudy (jdoe0001):

http://mathb.in/12801

OpenStudy (anonymous):

umm ok?

OpenStudy (jdoe0001):

so... what would you get for the right-hand side?

OpenStudy (anonymous):

-2??

OpenStudy (jdoe0001):

-2? how so?

OpenStudy (anonymous):

well 36 = 6^2 and its 1/36 so 6^-2?

OpenStudy (jdoe0001):

http://mathb.in/12802

OpenStudy (anonymous):

can you just help instead of the stupid link

OpenStudy (jdoe0001):

I'd do it right here, sure sadly the tools at this website are down, they've been down for about 3 days now, and is still unresolved, thus the link shows the procedure, so, as you know a logarithm is a function returning an exponent

OpenStudy (anonymous):

then type it out

OpenStudy (jdoe0001):

depending on the material, that works, or may not work

OpenStudy (anonymous):

well just try I don't wanna use links im running slow as it is

OpenStudy (anonymous):

Don't you have a book or something to read in ?

OpenStudy (anonymous):

no if I did I wouldn't be here

OpenStudy (anonymous):

http://prntscr.com/2lpicx here's some really basic rules, i scanned my booj about exponents

OpenStudy (anonymous):

it's the top part